OpenStudy (anonymous):

can someone please help me with this problem -4x+y=-4 -2x+4y=-16 solve each system by elimination thanks anybody

jimthompson5910 (jim_thompson5910):

-4x+y=-4 -2x+4y=-16 -4x+y=-4 -2(-2x+4y)=-2(-16) -4x+y=-4 4x-8y=32 -7y = 28 y = 28/(-7) y = -4 ------------------------ -4x+y=-4 -4x-4=-4 -4x=-4+4 -4x=0 x=0/(-4) x=0 So the solutions are x = 0 and y = -4

OpenStudy (anonymous):

how did we get here: -2(-2x+4y)=-2(-16)

jimthompson5910 (jim_thompson5910):

I multiplied both sides by -2 because I wanted the coefficient of the x terms to add up to 0 and cancel out So in equation 1, I have -4x In equation 2, I have -2x, but I want 4x. So I must multiply that (and everything else in equation 2) by -2 to get that -2x to +4x

jimthompson5910 (jim_thompson5910):

from there, I added the two equations, the x terms then cancel out: -4x+4x = 0x = 0, which goes away